OpenStudy (anonymous):

What is the area of the triangle? Right triangle with base labelled twenty-four centimeters height labeled 7 centimeters and third side labeled twenty-five centimeters.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[A = \frac{ 1 }{ 2 }bh\] So your area is equal to half of the height times base. So the base is 24 and the height is 7. 24 * 7 = 168 *1/2 (also equal to .5 or 2) is 84. So the area is 84 cm^2

OpenStudy (anonymous):

So 7.24*7+168*1/2 is 84 So that's how you do it

OpenStudy (anonymous):

I also have another question, The rear window of each side of a car is in the shape of a triangle. If that piece of glass has area twenty-four square inches and height 8 inches, what is the base of that triangular glass?

OpenStudy (anonymous):

Not exactly. \[(7 \times 24) \div 2\] is what I did for the triangle

OpenStudy (anonymous):

Would it be 24/8=3 so the base of the triangular glass is 3

OpenStudy (anonymous):

Hmm hang on

OpenStudy (anonymous):

What is the area of the parallelogram? 

OpenStudy (anonymous):

I believe b = 6 actually

OpenStudy (anonymous):

\[A = \frac{ 1 }{ 2}bh\]\[24 = \frac{ 1 }{ 2 }8b\]\[24\times2 = 8b \]I got rid of the 1/2 by multiplying both sides by 2. So 48 = 8b, divide by 8, b = 6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

What about the paraelgram?

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@brandi2013

OpenStudy (anonymous):

is it 42??

OpenStudy (anonymous):

@TSwizzle

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

It's just 10*5...but why'd you ask me?

OpenStudy (anonymous):

I was asking random people but THANKS my cousin used to be on here and she new you

OpenStudy (anonymous):

what was her username?

OpenStudy (anonymous):

Eh, I believe teresa1369

OpenStudy (anonymous):

...I don't know her.

OpenStudy (anonymous):

oh she has been gone for a while though

OpenStudy (anonymous):

Oh well...sorry, I don't remember her.